# MAVFlow: Preserving Paralinguistic Elements with Conditional Flow Matching for Zero-Shot AV2AV Multilingual Translation

### v1

@@ 5.3 Evaluation
We assess our model by using speaker similarity metrics, SS (speaker similarity) leveraging ERes2Net [11] providing a robust measure of how closely the synthesized speech matches the target speaker’s identity. ERes2Net is a widely used model trained on the VoxCeleb2 [14] dataset for speaker classification. We also consider Mel Cepstral Distortion-Dynamic Time Warping (MCD-DTW) [9] and its speech-length weighted variant, MCD-DTW-SL [9], to evaluate synthesis quality. MCD-DTW assesses alignment between generated and target mel-spectrograms, while MCD-DTW-SL additionally incorporates speech duration, providing a more comprehensive quality metric. We then examine translation quality with the ASR-BLEU score. Specifically, an automatic speech recognition (ASR) system is used to transcribe the generated audio, and the resulting text is compared against the ground-truth transcription to calculate the BLEU score [47]. Additionally, to evaluate the accuracy of emotion recognition, we assessed the audio generated by each system using pretrained emotion2vec [39].
@@
For visual quality assessment, we employ Lip Sync Error (LSE) confidence and distance (-C/-D) [50] and Fréchet Inception Distance (FID) [21], where the LSE metrics quantify the synchronization accuracy of lip movements relative to the audio, while FID measures the distributional similarity between generated frames and real images. Additionally, to measure emotional accuracy from the generated visual frames, we utilize a 5 classes ***Neutral, Happy, Sad, Surprise, and Fear pretrained EmoFAN [58] model for emotion classification.
Table 1: Comparison of zero-shot speaker similarity scores between X-En translated speech and native speech for traditional cascaded systems and direct systems. En: English, Es: Spanish, Fr: French, It: Italian, Pt: Portuguese.
@@
Table 2: Comparison of zero-shot speaker similarity scores for traditional cascaded systems and direct systems, with additional evaluation on the CREMA-D dataset for emotion evalution.
@@ 5.4 Zero-shot Audio Translation Result
To evaluate how accurately the emotion in the speech generated after translation reflects the emotion of the original speech, we compare the proposed model with the baseline model (AV2AV) using the CREMA-D dataset. The evaluation is based on the emotional accuracy calculated by the emo2vec model, which examines how the target speech (the synthesized speech after translation) is classified into ground-truth emotion categories. In Table 2, the emo2vec [39] model achieves approximately 82% classification accuracy on the ground-truth audio, serving as an upper bound for the emotion recognition model itself. In this experiment, our model achieves 36.5% emotional accuracy (+2.8%, +7.91%, and +18.94% compared to AV2AV, ASR + YourTTS [7], and ASR + xTTS [8] respectively), suggesting that it successfully synthesizes speech that preserves emotional traits.
To focus on evaluating the zero-shot performance of AV-Render, all models used for evaluation were trained only on the LRS3 dataset, without any separate fine-tuning on the CREMA-D dataset. Therefore, the actual emotion recognition performance might be limited by the domain difference between datasets. For instance, LRS3 primarily consists of speech from English speakers in news and interview settings, which may differ in speaker style, tone, and pronunciation from CREMA-D, which involves more dramatic emotional expressions. Despite these discrepancies, MAVFlow shows improvements over AV2AV, demonstrating its ability in preserving not only emotion but also speaker characteristics (SS and DTW).
@@
AV2AV AV-AV 26.57 31.27 23.24 24.51
MAVFlow (ours) AV-AV 26.97 31.33 23.43 24.97
Translation Qualtiy.
@@
Table 6: Visual emotion recognition from generate visual frames on CREMA-D
Method Emo-Acc (%)
GT 31.95
AV2AV 30.61
MAVFlow (ours) 33.29
@@ 5.5 Zero-shot Video Translation Result
In Table 5, we evaluated the visual quality of the generated videos and the synchronization between the audio and visual components. As a result, MAVFlow achieved an LSE-C score of 8.30, outperforming all baseline methods. Particularly, when compared to AV2AV (C4), which has a similar direct synchronization structure to ours, MAVFlow demonstrated significant improvements across all metrics: LSE-C (+0.87), LSE-D (-0.49), and FID (-0.61). These results indicate that audio-visual guidance not only enhances the consistency of synthesized speech but also positively affects face generation quality.
@@
In Figure 3, we visually analyze the generated visual quality on the CREMA-D dataset and evaluate whether each generated visual frame accurately reflects the speaker’s emotion using a visual emotion recognition model. Through this evaluation, we confirm that our proposed method, which applies visual embedding at the frame level, effectively captures the original emotional state of the speaker over time. For instance, in Figure 3, the AV2AV method incorrectly predicted ’HAP’ (Happy) for an original video labeled with ’ANG’ (Anger). Upon examining each frame closely, it becomes clear that the video generated by AV2AV fails to adequately express the anger emotion, particularly around the mouth shape and nose area, compared to the ground truth video. These visual observations are further supported by the results in Table 6, which show that while AV2AV’s visual emotion recognition performance decreases compared to the ground truth, our proposed method demonstrates improved performance.
Ablation Study for Audio-Visual Guidance.
We conducted an ablation study to examine the effect of each modality guidance on audio generation. Table 7 presents the results of evaluating the effect of audio and visual modality guidance on emotion recognition using the CREMA-D audio dataset. Additionally, it includes the analysis of results from translating audio in Es, Fr, It, and Pt to En using the MuAViC dataset, based on the settings outlined in Table 1. The SS and MCD-DTW values in Table 7 were averaged across each language for analysis. As a result, we observed that when both audio and visual guidance were provided, speaker similarity and emotional accuracy improved. One interesting observation is that when visual guidance is provided alone, speaker similarity slightly increases or is maintained(as seen in Tables 7), but Emo-Acc decreases. This suggests that visual guidance alone has a minimal effect on maintaining emotion, and its complementary effect is maximized when combined with audio guidance.
@@ 6 Conclusion
In this paper, we introduced MAVFlow, a zero-shot audio-visual translation framework utilizing conditional flow matching to address speaker consistency challenges inherent in existing AV2AV methods. By effectively integrating paralinguistic characteristics from both audio and visual modalities, MAVFlow significantly enhances speaker consistency across languages without intermediate text representations. Our method leverages discrete speech units and dual-modal guidance to synthesize high-quality mel-spectrograms, resulting in improved lip synchronization, emotional accuracy, and overall visual quality. Experimental evaluations on the MuAViC and CREMA-D datasets confirm that MAVFlow outperforms prior AV2AV methods, establishing it as a robust and efficient solution for multilingual audio-visual translation.

### v2

@@ 5.3 Evaluation
We assess our model by using speaker similarity metrics. SS (speaker similarity) leverages ERes2Net [11], providing a robust measure of how closely the synthesized speech matches the target speaker’s identity. ERes2Net is a widely used model trained on the VoxCeleb2 [14] dataset for speaker classification. Since SS alone is insufficient to evaluate the temporal alignment between generated and target mel-spectrograms, we additionally adopt Mel Cepstral Distortion with Dynamic Time Warping (MCD-DTW) [9] and its speech-length weighted variant (MCD-DTW-SL) [9]. The SL variant further accounts for speech duration, providing a more comprehensive quality metric. We then examine translation quality with the ASR-BLEU score. Specifically, an ASR system is used to transcribe the generated audio, and the resulting text is compared against the ground-truth transcription to calculate the BLEU score [48]. Additionally, to evaluate the accuracy of emotion recognition, we assess the audio generated by each system using the pretrained emotion2vec [40].
@@
For visual quality assessment, we employ Lip Sync Error (LSE) confidence and distance (-C/-D) [51] and Fréchet Inception Distance (FID) [22], where the LSE metrics quantify the synchronization accuracy of lip movements relative to the audio, while FID measures the distributional similarity between generated frames and real images. Additionally, to measure emotional accuracy from the generated visual frames, we utilize a 6-class [^1] Neutral, Happy, Sad, Angry, Disgust, and Fear pretrained MAE-DFER [59] model for emotion classification. Also, emotion embedding cosine similarity (ES) is used to complement class-wise accuracy, which may miss subtle emotional variations due to its fixed set of classes.
Table 1: Comparison of zero-shot speaker similarity scores between X-En translated speech and native speech for traditional cascaded systems and direct textless systems. En: English, Es: Spanish, Fr: French, It: Italian, Pt: Portuguese.
@@
Table 2: Comparison of zero-shot speaker similarity scores of generated audio for traditional cascaded systems and direct systems, with additional emotion evaluation on the CREMA-D dataset.
@@ 5.4 Zero-shot Audio Translation Result
To evaluate how accurately the emotion in the speech generated after translation reflects the emotion of the original speech, we compare the proposed model with the baseline model (AV2AV) using the CREMA-D dataset. The evaluation is based on the emotional accuracy calculated by the emo2vec model, which examines how the target speech (the synthesized speech after translation) is classified into ground-truth emotion categories. In Table 2, the emotion2vec [40] model achieves approximately 82% classification accuracy on the ground-truth(GT) audio, serving as an upper bound for the emotion recognition model itself. In this experiment, our model achieves 36.5% emotional accuracy (+2.8%, +7.91%, and +18.94% compared to AV2AV, ASR + YourTTS [7], and ASR + XTTS [8] respectively), suggesting that it successfully synthesizes speech that preserves emotional traits.
@@
[…]
Translation quality.
@@
Table 6: Visual emotion recognition accuracy (Emo-Acc) and emotion embedding cosine similarity (ES) measured from the generated visual results on CREMA-D.
Method Emo-Acc (%) $\uparrow$ ES $\uparrow$
GT 76.83 1.00
AV2AV 67.20 0.87
MAVFlow (ours) 72.68 0.92
Table 7: Audio emotion accuracy and embedding cosine similarity (ES) after additional training (+: additional training on CREMA-D).
Method Emo-Acc (%) $\uparrow$ ES $\uparrow$ SS $\uparrow$
AV2AV 33.66 0.84 0.33
MAVFlow 36.46 0.86 0.39
MAVFlow + 51.46 0.90 0.49
@@ 5.5 Zero-shot Video Translation Result
In Table 5, we evaluated the visual quality of the generated videos and the synchronization between the audio and visual components. MAVFlow achieves an LSE-C score of 8.30, outperforming all baseline methods. Particularly, when compared to AV2AV (C4), which has a similar direct synchronization structure to ours, MAVFlow demonstrates significant improvements across all metrics: LSE-C ($+$0.87), LSE-D (--0.49), and FID (--0.61). These results indicate that audio-visual guidance not only enhances the consistency of synthesized speech but also positively affects face generation quality.
@@
In Figure 3, we analyze the generated visual quality on the CREMA-D dataset and evaluate whether each generated visual frame accurately reflects the speaker’s emotion using a visual emotion recognition model. Through this evaluation, we confirm that our proposed method, which applies visual embedding at the frame level, effectively captures the original emotional state of the speaker over time. For instance, in Figure 3, the AV2AV method incorrectly predicted ‘HAP’ (Happy) for an original video labeled with ‘ANG’ (Anger). Upon examining each frame closely, it becomes clear that the video generated by AV2AV fails to adequately express the anger emotion, particularly around the mouth, compared to the ground truth video. These visual observations are further supported by the quantitative results in Table 6, which show that while AV2AV’s visual emotion recognition performance decreases compared to the ground truth, our proposed method demonstrates better preservation not only in terms of accuracy (Emo-Acc) but also in embedding similarity (ES). Additional visual quality can be referred to in Appendix A.
5.6 Ablation Study
Additional training on emotion dataset.
In Table 2, we only trained our model on the MuAViC dataset to enable zero-shot evaluation on the unseen CREMA-D benchmark. However, additional training on emotion-rich audio-visual datasets can significantly enhance emotion transfer performance. In Table 7, we lightly uptrained MAVFlow on CREMA-D training datasets (referred to as MAVFlow+), resulting in a notable increase in Emo-Acc from 36.46% to 51.46% as well as an improvement in ES from 0.86 to 0.90. Since MAVFlow outperformed the baselines using only MuAViC, we expect that incorporating such emotion-rich data would further widen this performance gap.
Effect of audio-visual guidance.
We conducted an ablation study to examine the effect of each modality guidance on audio generation. Table 8 presents the results of evaluating the effect of audio and visual modality guidance on emotion recognition using the CREMA-D audio dataset. Additionally, it includes the analysis of results from translating audio in Es, Fr, It, and Pt to En using the MuAViC dataset, based on the settings outlined in Table 1. The SS and MCD-DTW values in Table 8 were averaged across each language for analysis. As a result, we observed that when both audio and visual guidance were provided, speaker similarity and emotional accuracy improved. One interesting observation is that when visual guidance is provided alone, speaker similarity slightly increases or is maintained (as seen in Table 8), but Emo-Acc decreases. This suggests that visual guidance alone has a minimal effect on maintaining emotion, and its complementary effect is maximized when combined with audio guidance.
@@ 6 Conclusion
In this paper, we introduced MAVFlow, a zero-shot audio-visual translation framework utilizing Conditional Flow Matching (CFM) to address speaker consistency challenges inherent in existing AV2AV methods. By effectively integrating paralinguistic characteristics from both audio and visual modalities, MAVFlow significantly enhances speaker consistency across languages without intermediate text representations. Our method leverages discrete speech units and dual-modal guidance to synthesize high-quality mel-spectrograms, resulting in improved lip synchronization, emotional accuracy, and overall visual quality. Experimental evaluations on the MuAViC and CREMA-D datasets confirm that MAVFlow outperforms prior AV2AV methods, establishing it as a robust and efficient solution for multilingual audio-visual translation.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No. 2022-0-00641, XVoice: Multi-Modal Voice Meta Learning], [No. RS-2024-00457882, AI Research Hub Project], and [No. 2019-0-00075, Artificial Intelligence Graduate School Program (KAIST)].
Appendix
Appendix A Qualitative Results and Analysis
In Figure 4, we present dynamic emotional changes across frames within a single video at the first three frames from neutral to disgust. While MAVFlow effectively captures the emotional change of Ground Truth (GT) video from 15th frame, reflecting the shift starting from the 14th frame. AV2AV fails to reflect the emotion until around the 36th frame. Additionally, overall, MAVFlow better expresses emotions as well as the arousal level, which is indicated by the distance of a red dot from the center. Cascaded systems have been excluded from the comparison due to poor temporal alignment and their inability to embed emotional cues into the audio, which results in TFG output cannot reflect emotional expressions in the video. The DTW and DTW-SL metrics in Table 1 and Table 2, further confirm the notably poor temporal alignment of the cascaded systems.
Appendix B Class-Wise Emotional Analysis
B.1 Audio Emotional Results
In Table 9, we evaluate the class-wise emotion recognition accuracy of the generated audio using the pretrained emotion2vec [40]. Compared to AV2AV, MAVFlow shows slightly lower performance for the Sad, Disgust, and Fear classes, while demonstrating comparable or superior results for Happy, Neutral, and Angry. Notably, MAVFlow exhibits a significant advantage in the Angry class, ultimately achieving better overall performance than AV2AV in both Emo-Acc and ES metrics (as shown in Table 7). Furthermore, the MAVFlow + model, trained with additional emotional datasets, achieves improved performance across most emotion classes, with a substantial gain in overall Emo-Acc.
Table 9: Class-wise emotion accuracy (%) of generated audio (+: additional training on CREMA-D).
[…]
GT 89.29 85.00 89.17 89.29 77.86 62.14 81.95
AV2AV 30.00 22.86 80.00 28.57 30.71 16.43 33.66
MAVFlow 36.43 11.43 80.00 62.86 20.00 14.29 36.46
MAVFlow + 69.29 22.86 66.67 80.71 32.86 38.57 51.46
B.2 Visual Emotional Results
In Table 10, we evaluated class-wise visual emotion accuracy using pretrained MAE-DFER [59]. Also, follow MAE-DFER, we report both Unweighted Average Recall (UAR) and Weighted Average Recall (WAR) as evaluation metrics. UAR calculates the average recall by treating each class equally, which helps account for class imbalance, while WAR weights the recall by the number of samples per class, reflecting the actual class distribution in the dataset. As a result, MAVFlow achieved strong performance in terms of both UAR and WAR, particularly excelling in the angry, disgust, and fear emotion classes.
Table 10: Class-wise emotion accuracy, unweighted and weighted average recall (UAR%, WAR%), and ES of the generated visuals, all measured with MAE-DFER (+: additional training on CREMA-D).
Method Happy Sad Neutral Angry Disgust Fear UAR WAR ES
GT 97.14 67.86 76.67 78.57 87.86 52.86 76.83 76.83 1.00
ASR+YourTTS+TFG 89.86 60.71 72.88 50.71 83.57 40.00 66.29 66.05 0.85
ASR+XTTS+TFG 94.93 55.00 72.03 72.86 85.71 31.43 68.66 68.50 0.91
AV2AV 95.00 64.29 79.17 62.14 77.14 27.14 67.48 67.20 0.87
MAVFlow 95.00 53.57 75.00 80.71 88.57 43.57 72.74 72.68 0.92
MAVFlow + 95.00 63.57 78.33 76.43 87.14 37.86 73.06 72.93 0.93
Appendix C Inference Time Comparison
MAVFlow does not rely on intermediate text representations, resulting in faster inference compared to the cascaded system. Furthermore, it is more efficient by applying the speed-friendly CFM module compare to diffusion model. We compared the inference speed using one A6000 GPU, observing processing times of 1.66s for MAVFlow, 1.22s for AV2AV, and 1.75s for the 4-cascaded model to handle a 2.35s audio-visual input through the complete pipeline.
Appendix D Limitation
MAVFlow currently leverages emotional embeddings only from face and speaker embeddings from audio. However, we believe that incorporating emotional cues from audio (e.g., prosody, timbre, and other paralinguistic features) into the guidance of CFM could further enhance performance. Furthermore, since we directly adopt the unit extractor and unit-to-unit translation modules from previous work [13], improving semantic translation quality remains an open challenge.
Figure 4: Additional qualitative comparison for frame-level analysis. Each row shows GT, MAVFlow, AV2AV, and Cascade (ASR+XTTS+TFG), respectively.